OpenStudy (anonymous):

what is the perimiter if w= square root of 13 and L= 7 square root of 13 a)91 b)8 square root 13 c) 14 square root 13 d) 16 square root 13

OpenStudy (anonymous):

Okay, so first you need to know the rules of multiplying square roots. They basically go as follows: \[2\sqrt{7}*\sqrt{7}=2\sqrt{7*7}=2\sqrt{49}\] --> Basically you just multiply the radicand (number inside the square root) with the other one. So in your problem, the equation would look like this: \[\sqrt{13}*7\sqrt{13}=7\sqrt{13*13}=7\sqrt{169}\] so the final answer is sqrt{169}